Hi, I have a question. How would you solve this equation..
-(x+1)=2(3x-1)

Answer by stanbon(75887)

-(x+1)=2(3x-1)
-----
-x - 1 = 6x -2
7x = 1
x = 1/7
===========
Cheers,
Stan H.
===========